Assuming a normal distribution with a true mean of 80.3 Pascals and a standard deviation of 2.3 Pascals, what is the probability (in percentage) that future measurements will fall above 83.5 Pascals?
Answer #1

Here, X ~ Normal(mean = 80.3, sd = 2.3).
We know, (X - 80.3)/2.3 = Z ~ Normal(mean = 0, sd = 1).

Now, Prob.(X > 83.5) = P(Z > (83.5 - 80.3)/2.3) = P(Z > 1.3913) = 0.0821 = 8.21%.